### v3

@@ 5.1 Average treatment effects
In Figure 2, we display the proportions of participants employed in technology jobs over the span of 16 months, post-application to the program. This is illustrated across all experimental groups. [^13] We construct outcomes over time based on the new job’s start date. In some cases, we only observe the year the new job started rather than the specific month. For those observations, we assign December of that year as the starting month. Consequently, the average outcomes at the 4th month after the end of either program are lower than the average of tech job 4 months across the experimental groups. In the left figure, we consider three main groups: participants of Mentoring, participants of Challenges, and all other applicants to Dare IT (we refer to this group as Out of Dare IT). We can see that all groups start at zero, with none of the applicants having a job in technology; over time, the share of subjects that have a tech job in each group increases, with the Mentoring group increasing right from the moment the program started, and achieving over 50% probability of having a job in tech after 8 months. In contrast, the Challenges group follows a similar trajectory to the Out of Dare IT group until the 8th month, which is approximately 3 months after the end of the program, but then the gap between those two groups increases to almost 20 percentage points.
The difference in average outcomes across the three groups appears to be substantial, particularly in the case of Mentoring. However, the subjects in the Mentoring group were selected by mentors; thus, the program’s impact on the probability of getting a tech job might be compounded by the influence of the non-random selection of the sample. In the right figure, we decompose the effect of selection from the impact of Mentoring. We show the outcomes across applicants selected by mentors and randomized into Mentoring, selected, but randomized into control, and those that were not selected (the non-selected group consists of the non-selected applicants to Mentoring, and the control group, and late applicants to Challenges). We observe that in the 16th month, the probability of having a job in technology differs by over 20 percentage points between selected and not-treated and non-selected applicants, and it is substantially larger than the difference between the two groups that were selected but randomized into treatment or control. On average, mentors selected candidates that were more likely to get a job in technology. This observation highlights that, in this context, the ability to carry out a randomized trial is crucial for obtaining reliable estimates of the program’s effectiveness.
Figure 2: Mean outcome per group over time.
@@
In Table 2, we present the estimates of the average treatment effects for both the Mentoring and Challenges programs. Our analysis covers both short-term and longer-term outcomes: securing a job in tech within 4 months and 12 months after the programs, respectively. [^14] Tech job 4 months refers to 4 months after the end of a program which corresponds to approximately 8 months after application; analogously for Tech job 12 months. Additionally, we assess the influence of the programs on the acquisition of any new job within 4 months following the end of a program. For the short-term outcomes, we employ the difference-in-means estimator. [^15] Given the paired design of the mentoring experiment, we estimate variance as $\hat{V}(\hat{\tau}^{M})=\frac{1}{J\times(J-1)}\sum_{g=1}^{J}(\hat{\tau}^{M}_{j}-\hat{\tau}^{M})^{2}.$ (2) Athey and Imbens (2017) show that this is a conservative estimator of variance in a paired experiment. In the case of Challenges, we use the Splawa-Neyman estimator of variance. When estimating the longer-term effect, we utilize the fact that we observe monthly outcomes to estimate a Cox proportional hazard model (Cox, 1972). In Appendix C, we provide an estimate for short-term outcomes using the Augmented Inverse Propensity Weighting estimator (Robins et al., 1994; Athey et al., 2019). Furthermore, Appendix D contains alternative specifications of the Cox proportional hazard model, and the results align with those shown in Table 2. The test results of proportionality of hazard rates can be found in Appendix G.
Table 2: Average treatment effect estimates.
@@
Difference-in-means Proportional hazard Difference-in-means Proportional hazard
New job Tech job 4 months Tech job 12 months New job Tech job 4 months Tech job 12 months
ATE 0.045 0.129 0.126 0.040 0.089 0.071
(0.05) (0.05) (0.025) (0.05) (0.04) (0.02)
ATE % baseline 10.749 43.941 36.887 13.636 45.306 46.972
(14.23) (17.09) (7.34) (12.00) (14.07) (10.01)
No. of obs. treatment 152 152 2265 183 183 2490
No. of obs. control 147 147 2205 225 225 3330
@@
Note: The estimates of the average treatment effects of Mentoring and Challenges on the probability of finding a new job, a new job in technology within 4 months, and within 12 months. The first row shows estimates of the average treatment effect using the difference-in-means estimator in the case of new job and tech job 4 months (see Appendix C for AIPW estimates). The average treatment effect on Tech job 12 months is estimated using Cox proportional hazard model (we present estimates from a model with no covariates, see Appendix D for Cox proportional models with covariates and with time-varrying treatment effects). The second row presents the estimates from the first row divided by the mean outcome in the control group. Rows three and four present the number of observations per experimental group. The first, second, and third columns present the results of Mentoring intervention. Columns four, five, and six show the results from the Challenges program. Standard errors are presented below the average treatment effect estimates. The variance of the difference-in-means estimates of the average treatment effect of Mentoring are estimated using Equation 2. The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator. Standard errors of the estimates from Cox proportional hazard model are estimated using the partial likelihood function (coxph implementation in R).
We find that participation in the Mentoring program significantly increases the probability of getting a job in the tech industry, by 13 percentage points within both 4 and 12 months after the program’s conclusion. The short-term effect’s estimate has a standard error of 5 percentage points, whereas the longer-term effect’s standard error is reduced to 2.5 percentage points, thanks to the proportional hazard model’s functional form assumption and our time-series data. The Challenges program also significantly enhances the likelihood of securing a tech job, although its impact is somewhat less pronounced than that of the Mentoring program. We estimate that participation in Challenges increases the chance of securing a tech job by 8.9 percentage points within 4 months (with a standard error of 4), and by 7.1 percentage points within 12 months (with a standard error of 2).
In the second row of Table 2, we express the average treatment effects as a percentage of the control group’s mean outcome. As previously noted, mentor selection leads to higher average outcomes in the Mentoring control group compared to the Challenges control group. As a result, when expressed in percentage terms, the impact of both programs appears quite similar. Participation in Mentoring increases the odds of securing a tech job within 12 months by 37%, whereas in Challenges by 47%. Importantly, we find no evidence of the treatment effect diminishing over time when comparing the programs’ effectiveness 4 and 12 months after their conclusion.
Finally, columns 1 and 4 of Table 2 display the estimated impact of participating in Dare IT programs on the probability of acquiring any new job within 4 months post-graduation. We do not observe any statistically significant difference in this outcome between treatment and control. Therefore, our results suggest that the treatment and control groups of both experiments managed to secure new jobs at similar rates. However, those in the treatment groups were more likely to find employment in the technology sector.
@@
In Appendix F, we show the results based on the outcomes survey. We find that the survey respondents in both treatment groups report a higher number of job offers and higher salaries than the control group. The differences between treatment and control are greater in the case of Mentoring than Challenges.
@@ 5.2 Heterogeneous treatment effects
We explore heterogeneity in treatment effects by conducting an analysis of treatment effects across different subgroups using the Cox proportional hazard model. We categorize applicants to the Mentoring and Challenges programs into non-overlapping groups based on four different types of characteristics: STEM education, length of professional experience, type of residence, and age. For each of these categories, we estimate a hazard model that includes treatment and subgroup indicators along with their interaction. The results of this analysis are presented in Table 3. It’s important to note that applicants in each of the non-overlapping groups may have other varying characteristics; to account for this, we include a full set of controls and interactions in Appendix I in Table 10. Furthermore, here we present only selected subgroups; Appendix I presents all the interaction terms.
@@
Baseline CATE Baseline CATE Diff. Mentoring - Challenges
@@
Less than 5 years 0.429 (0.03) 0.008 (0.04) 0.119 (0.01) 0.136 (0.03) -0.128 (0.05)
More than 5 years 0.344 (0.02) 0.142 (0.03) 0.178 (0.01) 0.021 (0.02) 0.121 (0.03)
Difference subgroups -0.085 (0.03) 0.134 (0.05) 0.059 (0.02) -0.115 (0.03) 0.248 (0.06)
Type of residence
Small town 0.307 (0.03) 0.285 (0.05) 0.161 (0.02) 0.160 (0.03) 0.126 (0.06)
Warsaw 0.353 (0.02) 0.075 (0.03) 0.148 (0.01) 0.051 (0.02) 0.023 (0.03)
Difference subgroups -0.046 (0.03) 0.211 (0.05) 0.013 (0.02) 0.109 (0.04) 0.102 (0.07)
Age
Over 30 years old 0.261 (0.02) 0.125 (0.03) 0.130 (0.01) 0.114 (0.02) 0.011 (0.04)
Under 30 years old 0.428 (0.02) 0.117 (0.04) 0.178 (0.01) 0.025 (0.02) 0.092 (0.04)
Difference subgroups -0.167 (0.03) 0.008 (0.05) -0.049 (0.02) 0.089 (0.03) 0.011 (0.05)
Note: Estimates of average treatment effects conditioned on subgroups in Mentoring and Challenges. The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazard models, which include treatment indicator, subgroup indicator, and interaction. The first and third columns show the subgroup’s average outcome within control subjects. The second and fourth columns show the difference between treatment and control in the subgroup. The last column shows the difference between Mentoring and Challenges of the treatment effects. Standard errors are in parentheses.
First, we observe that participants with a STEM degree have higher baseline outcomes than those without a STEM degree. Both programs, Mentoring and Challenges, are more effective for participants without STEM degrees, bridging the outcome gap across these two groups. Specifically, in the case of Mentoring, we find that the treatment entirely eliminates the STEM outcomes gap. When comparing the two programs, we find no statistically significant difference for participants with STEM degrees, while for those without STEM degrees, Mentoring leads to nearly 15 percentage points higher treatment effects than Challenges. Second, when focusing on the subgroups with more or less than 5 years of professional experience, we find that Challenges is more effective for participants starting their professional career, with nearly 13 percentage points higher treatment effects in this subgroup. Conversely, for participants with longer professional experiences, Mentoring appears more effective. Third, we find that participants from smaller towns benefit more from both programs. However, this difference is particularly pronounced in the case of Mentoring. Participants from smaller towns have lower baselines than participants from Warsaw, but after the treatment, the gap is entirely closed. In fact, Mentoring participants from smaller towns achieve, on average, higher outcomes than Mentoring participants from Warsaw. Finally, participants below 30 years old have higher baselines than those over 30 years old. However, both programs help close the gap between the two age groups by having higher treatment effects for older participants.
@@
In the case of Mentoring, we observe several additional characteristics; these come from the Mentoring registration survey, which was more extensive than the Challenges one, and we also observe from characteristics of mentors. In Table 4, we present additional results documenting heterogeneity in treatment effects along those additional dimensions.
@@
Variable value is 1 Variable value is 0
@@
Mother 0.348 (0.02) 0.288 (0.04) 0.398 (0.02) 0.068 (0.03) 0.220 (0.05)
Applied before 0.384 (0.03) 0.141 (0.05) 0.330 (0.02) 0.121 (0.03) 0.020 (0.05)
@@
Managerial exp. 0.285 (0.02) 0.223 (0.03) 0.414 (0.02) 0.012 (0.03) 0.211 (0.05)
Long exp. tech 0.288 (0.02) 0.226 (0.04) 0.389 (0.02) 0.052 (0.04) 0.174 (0.05)
1st time mentor 0.354 (0.02) 0.131 (0.03) 0.342 (0.02) 0.112 (0.03) 0.019 (0.05)
@@
Note: Estimates of average treatment effects conditioned on subgroups in Mentoring. The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazard models, which include treatment indicator, subgroup indicator, and interaction. Applied before takes the value of 1 if the mentee has applied to the program in the past, but did not get selected into the program, and reapplied, and the value of zero if this is the first time applying to the program. Long exp. in tech takes the value of 1 if the mentor worked for over 5 years in the technology sector (the median value), and zero otherwise; 1st time mentor takes the value of 1 if this is the first time the mentor provides formalized mentoring service, and zero otherwise. Columns 2 and 3 describe the values for subgroups where the subgroup indicator takes the value 1, while columns 4 and 5 when the indicator takes the value of 0. The last column is the difference in the effectiveness of Mentoring across the two subgroups. Columns 1 and 3 are average outcomes in control groups, and columns 2 and 4 are the estimates of the average treatment effect. Standard errors are in parentheses.
First, we observe that the Mentoring program is particularly effective for mothers. Mothers who were selected by mentors but randomized into the control group had a baseline probability of 35% of finding a job in technology, compared to 40% for those without children. However, the average treatment effect for mothers is significantly larger at 29 percentage points, compared to just 7 percentage points for those without children. As a result, Mentoring not only enables mothers to catch up with applicants without children but even surpass them in the probability of getting a job in technology. In contrast, we do not observe a high difference in the effectiveness of Mentoring between first-time applicants and those who applied in previous editions.
Turning our attention to the characteristics of mentors, we find that mentors with managerial experience and those with a long history in the technology sector (over 5 years) tend to select, on average, less promising candidates, as indicated by baseline outcome probabilities of 29% for both groups, compared to the baseline of 41% for mentors without managerial experience and 39% for those with short tenure in tech. However, these mentors with managerial experience and long tech sector tenure are highly effective, with their mentees achieving a probability of over 50% of securing a job in technology. Interestingly, we also find that first-time mentors are as effective as mentors who have had previous mentoring experience. These findings are consistent with the notion that effective mentors leverage their knowledge, skills, and relationships accumulated over years of working in the technology sector, rather than improving their mentoring capabilities through the practice of mentoring.
Patterns of treatment effects heterogeneity are consistent with some of the mechanisms through which the two programs might impact employability that we hypothesized in Section 3. Participants from smaller towns and those without family or friends working in tech (see Table 10) might have less extensive professional networks in the technology sector, and mentors with longer experience in tech are likely to have richer networks. Thus, the finding that participants with these characteristics benefit more from Mentoring is consistent with the mechanisms that Mentoring helps mentees by bolstering their professional network and enhancing their capacity to utilize these networks effectively. As for Challenges, we hypothesized that the program would equip participants with credible signals of their technical skills. This mechanism is consistent with the finding that participants without STEM degrees and those with shorter professional experience, thus fewer other signals on their resumes, exhibit larger benefits.

### v4

@@ 5.1 Average Treatment Effects
In Figure 2, we display the proportions of participants employed in technology jobs over the span of thirteen months post-application to the program. This is illustrated across all experimental groups. [^8] We construct outcomes over time based on the new job’s start date. In some cases, we only observe the year the new job started rather than the specific month. For those observations, we assign December of that year as the starting month. across the experimental groups. In the left figure, we consider three main groups: participants of Mentoring, participants of Challenges, and all other applicants to Dare IT (we refer to this group as Out of Dare IT). We can see that all groups start at zero, with none of the applicants having a job in technology; over time, the share of subjects that have a tech job in each group increases, with the Mentoring group increasing right from the moment the program started, and achieving over 50% probability of having a job in tech after 8 months. In contrast, the Challenges group follows a similar trajectory to the Out of Dare IT group until the 8th month, which is approximately 3 months after the end of the program, but then the gap between those two groups increases to almost 20 percentage points. This differential timing, with Mentoring effects emerging during the program while Challenges effects manifest several months post-completion, suggests distinct mechanisms. These trajectories are consistent with the distinction between immediate network-based job matching versus the creation of portfolio signals and gradual employer response to them.
The right panel decomposes selection from treatment effects in Mentoring. Applicants selected by mentors but randomized to control achieve substantially higher technology employment (43% at month thirteen) than non-selected applicants (19%), confirming that mentors identify promising candidates. Furthermore, the treatment-control gap among selected applicants is smaller than the selection effect itself, underscoring the importance of randomization for isolating program impacts from mentor selection.
Figure 2: Share of Applicants with a Tech Job per Group Over Time.
@@
Table 2 presents estimated average treatment effects. We estimate the averate treatment effects using the difference-in-means estimator (columns 1-3, 5-7) and using Cox proportional hazard models that exploit monthly outcome variation (columns 4, 8). Appendix F verifies the proportional hazards assumption; Appendix C presents Kaplan-Meier survival curves confirming the timing patterns in Figure 2.
Both programs generate substantial, statistically significant increases in technology employment. Mentoring increases the probability of technology employment by 15.3 percentage points (SE 5.7), representing a 34% increase relative to the control group baseline of 45.6%. Challenges produces an 11.3 percentage point effect (SE 4.8), a 42% increase relative to the 26.6% control baseline. Both estimates use difference-in-means; results are nearly identical using Cox proportional hazard models that exploit monthly variation in job start dates (14.5pp for Mentoring, 11.2pp for Challenges).
Table 2: Average Treatment Effects of Mentoring and Challenges.
Difference-in-means Prop. hazard Difference-in-means Prop. hazard
@@
New job Tech job Non-tech job Tech job Cox New job Tech job Non-tech job Tech job Cox
ATE 0.078 0.153*** -0.076** 0.145*** 0.072 0.113** -0.041 0.112**
SE (0.057) (0.057) (0.031) (0.055) (0.051) (0.048) (0.042) (0.051)
ATE % baseline 13.576 33.676** -65.641** 31.718*** 14.458 42.646** -17.600 42.232**
SE (9.893) (14.783) (26.720) (11.992) (10.268) (18.131) (17.978) (19.301)
No. obs. treated 151 151 151 2265 166 166 166 2656
No. obs. control 147 147 147 2205 218 218 218 3488
@@
Note: The estimates of the average treatment effects of Mentoring and Challenges on the probability of finding a new job, a new job in technology, and a new job not in technology. Columns 1-3 and 5-7 present estimates using the difference-in-means estimator and columns 4 and 8 using a proportional hazard model with a full set of covariates. The first row shows estimates of the average treatment effect in the difference in the probability of reporting employment at month 15. The third row presents the estimates from the first row divided by the mean outcome in the control group. Standard errors are presented below the average treatment effect estimates. The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator. *** p<0.01, ** p<0.05, * p<0.1.
Job Search Selectivity.
Estimates in Column 3 of Table 2 suggest that Mentoring operates partly through increased job search selectivity. While Mentoring increases overall employment by only 7.8 percentage points (SE 5.7), statistically insignificant, it increases technology employment by 15.3 percentage points and decreases non-technology employment by 7.6 percentage points (SE 3.1). This implies that approximately half of Mentoring’s impact reflects participants becoming more selective, holding out for technology positions rather than accepting readily available employment in other sectors. In contrast, Challenges shows no significant effect on non-technology employment (-4.1pp, SE 4.2), consistent with this program operating through skill signaling rather than changing search behavior or persistence. The differential patterns support distinct mechanisms: Mentoring increases both participants’ capability to access technology jobs (through networks and tacit knowledge) and their willingness to persist in pursuing them (for example, through confidence-building), while Challenges primarily improves application success conditional on applying. This finding provides evidence consistent with the self-efficacy channel hypothesized in Section 3: mentors help women maintain focus on technology careers.
@@
In Appendix E, we show the results based on the outcomes survey. We find that the survey respondents in both treatment groups report a higher number of job offers and higher salaries than the control group. The differences between treatment and control are greater in the case of Mentoring than Challenges.
5.2 Impact of Program Participation on the Number of LinkedIn Connections
We directly test the network mechanism by examining whether program participation increases professional network size, measured by LinkedIn connection counts. This analysis distinguishes network expansion from other channels: if Mentoring works solely through skill development or signaling, we should observe no differential effect on connections relative to Challenges. Conversely, if networking is a primary mechanism, Mentoring should increase connections while Challenges, which operates entirely online without mentor introductions or employer contact, should not.
The data support the network mechanism. Mentoring increases LinkedIn connections by 52.7 (SE 20.8) from a control group baseline of 225, a 23% increase. In contrast, Challenges produces a statistically insignificant reduction of 11.4 connections (SE 17.8) from a baseline of 232. The differential treatment effect of 64 connections between programs (SE 28) provides direct evidence that Mentoring expands professional networks while Challenges does not.
To understand where these effects appear in the connection distribution, Table 3 presents quantile regression estimates. Mentoring generates large, statistically significant increases at the 25th percentile (62 connections, SE 27) and median (92 connections, SE 35), but no detectable effect at the 75th percentile. This pattern indicates the program particularly benefits participants with initially sparse networks. In contrast, Challenges shows no significant effects at any quantile, confirming this program does not operate through networking channels.
Table 3: Treatment Effects on LinkedIn Professional Networks.
Quantile Treatment Effect Mentoring Treatment Effect Challenges Difference: Mentoring and Challenges
0.25 62** 5 57*
(27.159) (13.391) (30.281)
0.5 92*** 5 87*
(35.426) (29.470) (46.082)
0.75 5 -8 13
(47.004) (50.706) (69.141)
Note: The table provides the quantile regression results for the Challenges and Mentoring programs. The treatment effects of the experimental groups on LinkedIn connections are shown at the 25th, 50th, and 75th quantiles. The difference column captures the difference in the treatment effects between the Mentoring and Challenges programs. Estimates are obtained by regressing the treatment indicator on the number of connections. Standard errors are in parentheses. *** p<0.01, ** p<0.05, * p<0.1.
The magnitude of network expansion is economically meaningful. A 92-connection increase at the median represents substantial growth from the control baseline of 225, comparable to the network size differences between mid-career and senior professionals in our sample. Moreover, the increase occurs despite LinkedIn’s 500-connection display cap, which likely attenuates our estimates: 25% of Mentoring treatment participants hit this ceiling compared to 23% of controls, suggesting the true effect exceeds our measured 52.7-connection average.
This network expansion cannot be solely attributed to new employment. Both Mentoring and Challenges participants secure technology jobs at high rates, yet only Mentoring participants expand their networks. Had job acquisition been the primary driver of connection growth, through adding new colleagues and industry contacts post-employment, we should observe comparable increases in both programs. The absence of connection growth among Challenges participants, despite their 38% technology employment rate, indicates that Mentoring’s network effects operate independently of (and plausibly prior to) job acquisition. This timing is consistent with mentors actively facilitating introductions during the program, as reported in our qualitative interviews, rather than networks expanding passively after job placement.
@@ 5.3 Heterogeneous Treatment Effects
The hypothesized contrasting mechanisms underlying Mentoring and Challenges generate testable predictions about which participants benefit most from each program. We exploit treatment effect heterogeneity to provide evidence consistent with these mechanisms: if Mentoring operates primarily through network expansion (as suggested by LinkedIn connections analysis), effects should concentrate among participants with weak existing networks; if Challenges works through skill signaling, effects should be largest for those with weaker pre-treatment credentials. We estimate heterogeneous treatment effects using Cox proportional hazard models that estimate treatment effects in subgroups defined by the pre-treatment characteristics of interest. We control for the full covariate set.
Table 4 presents heterogeneous treatment effects across three participant characteristics that map to credential strength and network access. These patterns provide evidence on which populations each program serves most effectively and tests the distinct mechanisms we hypothesize.
The clearest differential pattern emerges from professional experience. Challenges generates 28 percentage points larger treatment effects for participants with under five years of experience compared to those with longer careers (SE 10), whereas Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced participants (SE 11). The cross-program differential is 49 percentage points (SE 11). This contrast directly supports our theoretical predictions: early-career women lack established work history to credibly signal their technical abilities, making the Challenges portfolio particularly valuable as a complement to sparse employment records.
Educational background reveals similar patterns. Recall that all participants were pre-screened for technical competency, so heterogeneity by educational background likely reflects differences in credential signals and professional networks, and to a lesser degree underlying ability. Among Mentoring participants, those without STEM degrees show 19 percentage points larger effects than those with STEM backgrounds (SE 11). This pattern is consistent with the network access: non-STEM participants followed educational paths outside technology-adjacent fields, limiting opportunities to build connections with peers entering the sector.
In contrast, Challenges shows no STEM heterogeneity (difference: 0.3pp, SE 11). This null result demonstrates that portfolio signals are equally valuable regardless of educational credentials—employers evaluate demonstrated capability rather than inferring skill from degrees.
Geography provides additional evidence on network mechanisms. Mentoring generates 26 percentage point treatment effects for participants from small towns (SE 10) compared to only 3 percentage points for Warsaw residents (difference: 23pp, SE 14, p=0.11). This pattern aligns with network barriers binding more tightly outside major urban centers, where local technology ecosystems are sparse and professional connections to the sector are limited. Mentor networks based in Warsaw and other major cities provide greater marginal value when participants lack local access to technology professionals. Challenges shows no significant geographic heterogeneity (difference: -2pp, SE 1.2), consistent with online portfolio signals functioning equally regardless of location, employers evaluate portfolios based on technical content rather than applicant geography.
@@
Baseline CATE Baseline CATE Diff. Mentoring – Challenges
@@
More than 5 years 0.412 (0.049) 0.202 (0.066) 0.31 (0.039) -0.003 (0.065) 0.205 (0.063)
Less than 5 years 0.556 (0.075) 0 (0.087) 0.19 (0.042) 0.281 (0.075) -0.281 (0.086)
Difference subgroups -0.144 (0.089) 0.203 (0.109) 0.120 (0.057) -0.284 (0.099) 0.486 (0.106)
Small Town
Small town 0.429 (0.085) 0.26 (0.101) 0.256 (0.067) 0.149 (0.13) 0.111 (0.108)
Warsaw 0.447 (0.082) 0.033 (0.103) 0.182 (0.044) 0.166 (0.089) -0.133 (0.093)
Difference subgroups -0.019 (0.118) 0.227 (0.144) 0.074 (0.157) -0.019 (0.118) 0.246 (0.197)
Note: Estimates of the average treatment effects conditioned on subgroups in Mentoring and Challenges. The outcome variable is Tech job, and estimates are obtained using Cox proportional hazards models per subgroup with a full set of covariates. The first and third columns show the average outcome per subgroup within the control groups. The second and fourth columns show the difference between the treatment and control groups in each subgroup. The last column shows the difference in the treatment effects between Mentoring and Challenges. Standard errors are in parentheses.
@@
The Mentoring registration survey was more extensive than the Challenges one, allowing us to observe more characteristics. We also obtain the characteristics of mentors from their registration survey. In Table 5, we present additional results documenting heterogeneity in treatment effects along these dimensions.
@@
Characteristic variable value is 0 Characteristic variable value is 1
@@
Mothers 0.237 (0.07) 0.348 (0.097) 0.532 (0.048) 0.071 (0.062) 0.277 (0.115)
Applied before 0.5 (0.087) 0.139 (0.109) 0.442 (0.047) 0.143 (0.06) -0.004 (0.124)
@@
Managerial exp. 0.378 (0.056) 0.259 (0.073) 0.549 (0.059) 0.002 (0.074) 0.257 (0.104)
Long exp. in tech 0.39 (0.063) 0.277 (0.085) 0.512 (0.054) 0.021 (0.066) 0.257 (0.107)
1st time mentor 0.444 (0.062) 0.155 (0.08) 0.476 (0.055) 0.1 (0.073) 0.056 (0.108)
@@
Note: Estimates of the average treatment effects conditioned on subgroups in the Mentoring program. The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazards models, which include the treatment indicator and the set of controls. We estimate the model per subgroup. "Long exp. in tech" takes the value of 1 if the mentor has worked for over 5 years in the technology sector (the median value). Columns 2 and 3 describe the values for subgroups when the subgroup indicator takes the value 0, while columns 4 and 5 describe the values when the indicator takes the value 1. The last column shows the difference in the effectiveness of Mentoring between the two subgroups. Columns 1 and 3 are the average outcomes in the control groups, and columns 2 and 4 are the estimates of the average treatment effect. Standard errors are in parentheses.
Mentors with managerial experience generate 26 percentage point larger treatment effects than those without (SE 10). Similarly, mentors with over five years of technology sector experience produce 26 percentage point larger effects (SE 11). These patterns are consistent with the network channel: managers and senior professionals possess richer networks, better access to hiring decision-makers, and greater ability to facilitate job placements.
Interestingly, mentor selection patterns reveal that experienced, managerial mentors choose less promising candidates. Control group participants matched to managerial mentors have 38% baseline technology employment compared to 55% for those matched to non-managerial mentors. Yet treated participants of managerial mentors achieve higher outcomes (64% vs. 64%), because the treatment effect is substantially larger. This suggests that experienced mentors either (i.) deliberately select candidates who need more help, recognizing their own capacity to assist difficult cases, or (ii.) value different candidate attributes than baseline employment probability.
Notably, first-time mentors are as effective as repeat mentors (difference: 6pp, SE 11). This finding indicates that mentoring effectiveness stems from mentors’ accumulated professional networks, technical knowledge, and industry standing—not from mentoring skill developed through practice. A senior engineer with extensive connections generates value on her first mentoring experience; repeated mentoring does not improve outcomes.
Mothers and Work-Family Barriers.
The largest heterogeneity we observe is for mothers. Among non-mothers, Mentoring generates 7 percentage point treatment effects (SE 6); among mothers, the effect is 35 percentage points (SE 10), a difference of 28 percentage points (SE 12). Mothers in the control group face substantial disadvantage, with 24% technology employment compared to 50% for non-mothers. Mentoring completely eliminates this gap: treated mothers achieve 59% employment, surpassing non-mothers.
This large effect for mothers is consistent with Mentoring addressing multiple barriers simultaneously. First, mothers may have weaker networks if career interruptions weakened professional connections. Second, work-family conflict concerns may reduce mothers’ confidence in pursuing demanding technology careers; mentor encouragement and role-modeling by successful mentors in tech may counter this.
This finding contributes to understanding gender gaps in technology beyond representation. Even among women who have acquired technical skills and are actively seeking technology employment, motherhood creates additional barriers. Traditional explanations emphasize mothers’ labor supply decisions or employer discrimination; our results suggest that confidence and network deficits also play important roles, and relatively light-touch mentoring interventions can substantially mitigate these barriers.
The contrasting heterogeneity across programs aligns with our theoretical framework. Network barriers might bind most tightly for geographically isolated women and mothers, the groups showing the largest Mentoring effects. Information frictions bind most for early-career women with limited work history, those showing the largest Challenges effects. The programs’ differential impacts across populations indicate that successful policies must target multiple, distinct barriers rather than assuming a single mechanism operates uniformly. Moreover, the patterns demonstrate that even within the selected population of women with baseline technical skills, substantial heterogeneity exists in which constraints bind most tightly, with implications for optimal program assignment that we explore in Section 6.